OpenStudy (anonymous):

Identify a factor of this trinomial 9x^2+3x-20

OpenStudy (matt101):

15*(-12) = -180 15+(-12) = 3 [(9x+15)(9x-12)]/9 = 0 (3x+5)(3x-4) = 0 x = -5/3, 4/3